Koroashk:

A survey determined that of the seventh grade students at Mills Middle School preferred round chocolate candies to hard candy. When Jennifer duplicated the survey, of the students changed their preference from chocolate candy to hard candy. What fraction of the students in the second survey preferred round chocolate candy?

mhanifa:

Can you share ss of this question, it seems incomplete

Koroashk:

ok

Koroashk:

mhanifa:

All you need to do is subtract the second fraction from the first one

mhanifa:

Can you do that?

Koroashk:

nvm I figured it out sorry I messed you guys I'll you when I need help

itsmehjay:

Be aware that they have the same denominator so there won't be any finding lcd's Also, the survey was duplicated, so 2x

IhelpUuHELPme:

The answer you put is incorrect. Round Chocolate Candy: \[\frac{ 5 }{ 19 }\] Hard Candy: \[\frac{ 3 }{ 19 }\] The value of the denominators(the value on the bottom) is the same, so all you need to do is subtract the numerators. \[5 - 3 = ?\]

itsmehjay:

@itsmehjay wrote:
Be aware that they have the same denominator so there won't be any finding lcd's Also, the survey was duplicated, so 2x
actually wait no nvm ;-;